OpenStudy (anonymous):

given slope m=-6,(8,9) what the equation of y=what ?

OpenStudy (anonymous):

Use the general form of an equation of a line:\[y - y_1 = m(x-x_1)\] where y_1 is your given value of y, and x_1 is your given value of x. Plug the three given values into the equation and solve for y, giving your equation.

OpenStudy (anonymous):

what do i do with -6 then?

OpenStudy (anonymous):

Plug -6 into where m is, so that you get \[y - 9 = (-6) • (x - 8)\] after plugging each given into the equation.

OpenStudy (anonymous):

my mind is fried from this pellet all day, so hen how do i multiply -6 with x-8? -6x-14?

OpenStudy (anonymous):

Kinda, you have the right concept but just need to carry out the multiplication more carefully. y - 9 = -6x + 48, which leaves y = -6x + 57.

OpenStudy (anonymous):

how did u get 57?

OpenStudy (anonymous):

Add 9 to the each side, so you'll get y - 9 + 9 = -6x + 48 + 9, on the left hand it simplifies to y, and on the right hand it simplifies to -6x + 57.

OpenStudy (anonymous):

\[2^{2}+3^{3}-30\times40 \]

OpenStudy (anonymous):

do i multiply the exponents?

OpenStudy (anonymous):

Erm..are you trying to simplify the expression into a value?

OpenStudy (anonymous):

yes i need to simplify it

OpenStudy (anonymous):

my mind id drawing a blank lol sorry to bother you if i am...i appreciate the help lol

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

Lol, don't worry about it. I don't really think there is an easy way of shortening it, but all you have to do is calculate 2^2 (which is 4), 3^3 (which is 27), and 30 x 40 (which is 1200). Add 4 to 3, and subtract 1200 to get your answer, which is -1,169 if I haven't made a silly mistake along the way. :P

OpenStudy (anonymous):

thank you thats what i was thinking was done but i just thought i may have done something wrong

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

how well are you in math?

OpenStudy (anonymous):

You're very welcome :) I'm not sure how good my skills are, but for reference I'm in AP calculus.